OpenStudy (anonymous):

\[\int_{2}^{3} \frac{u^3+1}{u^3-u^2} du\]

OpenStudy (anonymous):

\[1+ \frac{u^2-1}{u^3- u^2}\]

OpenStudy (anonymous):

\[\int_{2}^{3} \frac{u^3+1}{u^2(u-1)} du\]

OpenStudy (turingtest):

\[1+\frac{u^2+1}{u^3-u^2}\]I think

OpenStudy (anonymous):

yep it's plus 1

OpenStudy (anonymous):

\[\int_{2}^{3} 1 du +\int_{2}^{3}\frac{u^2+1}{u^3-u^2} du\]

OpenStudy (turingtest):

so we can do partial fractions\[1+\frac{u^2+1}{u^2(u-1)}=1+\frac Au+\frac B{u^2}+\frac C{u-1}\]

OpenStudy (anonymous):

can we leave the 1+ out while doing partial fractions?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

\[\frac{u^2+1}{u^2(u-1)}=\frac Au+\frac B{u^2}+\frac C{u-1}\]\[Au(u-1)+B(u-1)+Cu^2=u^2+1\]

OpenStudy (turingtest):

\[\frac{u^2+1}{u^2(u-1)}=\frac Au+\frac B{u^2}+\frac C{u-1}\]\[Au(u-1)+B(u-1)+Cu^2=u^2+1\]\[(A+C)u^2+(-A+B)u-B=u^2+1\]so\[B=-1\]which means\[A=-1\]which means\[C=2\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[-\frac {1}{u}-\frac {1}{u^2}+\frac 2{u-1}\] \[-\int_{2}^{3} \frac {1}{u}du-\int_{2}^{3} \frac {1}{u^2}du+\int_{2}^{3} \frac 2{u-1}du\] \[-ln| u| +\frac1u+2ln|u-1|\] evaluated from u=2 to u=3

OpenStudy (turingtest):

don't forget to pout the +1 back in before you integrate!

OpenStudy (turingtest):

put*

OpenStudy (anonymous):

oowww...yes thank you. I will pout that back in

OpenStudy (anonymous):

Thanks Turing!

OpenStudy (turingtest):

welcome!